A body with mass m = 100kg makes uniform rectilinear movements in which case it reaches the kinetic energy of Ek = 20 KJ. How fast was it in that case? V =?


1
Expert's answer
2021-01-16T17:24:37-0500

The definition of the kinetic energy is:


"E_k= \dfrac{mv^2}{2}"

where "m = 100kg" is the mass of the body, and "v" is its velocity. Since "E_k = 20kJ = 20\times 10^3J", obtain for velocity:


"v = \sqrt{\dfrac{2E_k}{m}}\\
v = \sqrt{\dfrac{2\cdot 20\times 10^3}{100}} = 20m/s"

Answer. 20 m/s.